OpenStudy (anonymous):

Please solve the equation √(5a+2)=-√(a+7)

OpenStudy (anonymous):

I got -3/2, please check my answer

OpenStudy (anonymous):

try squaring both sides 5a +2 = a +7 4a = 5 a = 5/4

OpenStudy (anonymous):

if you sub these two results into the original equation it won't be satisfied so there is NO solution. I f you draw two graphs of each side you will see they don't intersect

OpenStudy (anonymous):

the RHS is negative for all values of x. LHS cannot have negative values

OpenStudy (anonymous):

thnks jimmyrep i forgot the minus sign ....lol :)